OpenStudy (anonymous):

Derivative of x^2+4/3x help

OpenStudy (abb0t):

\(\large \frac{d}{dx} (a^n) = na^{n-1}\)

OpenStudy (anonymous):

its a fraction

terenzreignz (terenzreignz):

Coefficients don't matter that much :) @abb0t \[\large \color{red}* \frac{d}{dx} (a^n) = na^{n-1}\cdot \color{red}{\frac{da}{dx}}\]

terenzreignz (terenzreignz):

Just sayin' @abb0t :D

OpenStudy (anonymous):

I tried doing the quotation rule, but the answer the book gives is different than what I get

terenzreignz (terenzreignz):

quotient rule? A better way would be to treat \(\large \frac34 \) just like any constant.

OpenStudy (abb0t):

Sometimes the solution given in a book is simplified to make the solution look "pretty" however, there are also other ways to write a solution.

terenzreignz (terenzreignz):

If you have to get the derivative of \[\Large ax\] You wouldn't hesitate to say it's \(\large a\) (right?)

OpenStudy (anonymous):

\[\frac{ x ^{2}+4 }{ 3x }\]

OpenStudy (anonymous):

thats it

terenzreignz (terenzreignz):

Well, the case involving \[\large \frac34x\] isn't much different

terenzreignz (terenzreignz):

oh okay, that clears things up :D

terenzreignz (terenzreignz):

Quotient rule being \[\Large \frac{d}{dx}\left[\frac{f(x)}{g(x)}\right]=\frac{f'(x)g(x)-f(x)g'(x)}{[g(x)]^2}\]

OpenStudy (anonymous):

yes, thats what I used

OpenStudy (anonymous):

but I get a different answer than my book

terenzreignz (terenzreignz):

Okay, so it all boils down to knowing what f(x), g(x), f'(x), and g'(x) are

terenzreignz (terenzreignz):

What is your answer, by the way?

OpenStudy (anonymous):

I got \[\frac{ 3x ^{2}-12 }{ (3x)^{2} }\]

terenzreignz (terenzreignz):

Okay... and what does the book say?

OpenStudy (anonymous):

the book has this \[\frac{ (x-2)(x+2) }{ (3x)^{2} }\]

terenzreignz (terenzreignz):

hehe

OpenStudy (anonymous):

???? am I doing something wrong?

terenzreignz (terenzreignz):

okay, wellI'm pretty sure it's like this... \[\Large \frac{(x-2)(x+2)}{3x^2}\]

terenzreignz (terenzreignz):

no parentheses in the denominator

terenzreignz (terenzreignz):

or am I missing something?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

no thats what the book has

OpenStudy (anonymous):

so what am I doing wrong

terenzreignz (terenzreignz):

It's different that way. So... just keep calm and take deep breaths... \[\Large \frac{3x^2-12}{9x^2}\] Catch me so far?

OpenStudy (anonymous):

yea

terenzreignz (terenzreignz):

Okay, so why don't we... factor out a 3 in the numerator? \[\Large \frac{3(x^2-4)}{9x^2}\]

OpenStudy (anonymous):

ok, now i got it

terenzreignz (terenzreignz):

>:)

terenzreignz (terenzreignz):

So, everything all right now?

OpenStudy (anonymous):

i was be stupid. yep.

terenzreignz (terenzreignz):

Nah, it was the book being nitpicky :D